From the foot of a tower 90 m high a stone is thrown up so as to reach the top of the tower. Two seconds later another stone is dropped from the tower.The two stones meet at ?
 

Grade:11

2 Answers

Arun
5 years ago
Height of the building (d) = 90m
Let’s find the velocity of stone thrown up:
We know, v^2 – u^2= 2ad
Velocity of stone after reaching top of the building = 0
Acceleration is opposite to acceleration due to gravity = -9.8m/s^2
0- u^2= 2*(-9.8)*90
- u^2=-1764
Therefore, velocity of stone moving up= 42m/s
Distance at time t = h(t) =ut-1/2at^2
h(t) = 42t-4.9t^2 -----(1)
h1(T) = height of building – 1/2aT^2
T=t-2 (since second stone is thrown after two seconds)
H1= 90 – 4.9(t-2)^2
When the two stones meet, distance are same. So, Set h= h1
42t - 4.9t^2 = 90 -4.9t^2 +19.6t -19.6
42t-19.6t=90-19.6
22.4t = 70.4 => t = 3.14s 
Substitute the value of “t” in equation (1)
h =83.6m
So, the two stones meet at 83.6m 
 
 
 
Killer bot
2 years ago
Taati kha loo  moj karo padh leya karo fail hojo ge bencho phone rakh ke padh laa book toh lodeaa fail hogya tah rishtedar aake gaand marr dege  iss laiye padh or padh ne doo